OpenStudy (anonymous):

Find the product and simplify the result. (8a − 4b)2

OpenStudy (saifoo.khan):

\[(8a-4b)^2\] \[64 a^2-64 a b+16 b^2\]

OpenStudy (anonymous):

Actually its raised to the second power

OpenStudy (anonymous):

\[ 64 a^2-64 a b+16 b^2 \]

OpenStudy (anonymous):

48

OpenStudy (saifoo.khan):

from where in this world u get 48?

OpenStudy (anonymous):

lolz :D

OpenStudy (anonymous):

8 ^2 is 64 - 4^2 is 16 = 48

OpenStudy (anonymous):

Use this general formula, \[(a - b)^{2}= a^{2} - 2ab + b^{2}\] It should guve you the answers.

OpenStudy (saifoo.khan):

there are variables man!! wake up!!

OpenStudy (anonymous):

@Chicken :Nice approach!

OpenStudy (anonymous):

I@chiken_little:I am sure that there is no "could be" ;)

OpenStudy (saifoo.khan):

what i wrote.

OpenStudy (anonymous):

yes he is ...

OpenStudy (saifoo.khan):

Yes im.

OpenStudy (anonymous):

the asnwer is 64a^2-64ab+16b^2

OpenStudy (anonymous):

why 64^2 twice then wouldnt that apply to 16^2

OpenStudy (saifoo.khan):

No! man. there is a variable in between and that square is on top of a variable.

OpenStudy (anonymous):

I'm sorry you guys are all wrong, (8a - 4b) ^2 = \[64a^{2} - 32ab + 16b^{2}\]

OpenStudy (anonymous):

beg a differ, my good chum

OpenStudy (saifoo.khan):

o_O